OpenStudy (jordanloveangel):

What is the slope of the line passing through the points (2, 5) and (0, −4)?

OpenStudy (jordanloveangel):

@dan815

OpenStudy (jordanloveangel):

@undeadknight26 @KonaFirestar @fallenangelorchid

OpenStudy (anonymous):

i can't help, Jordan. Sorry.

OpenStudy (jordanloveangel):

@deadpoolroxz

OpenStudy (jordanloveangel):

its kk

OpenStudy (jordanloveangel):

@Kit_Kat_Nat_<3

OpenStudy (anonymous):

I honestly have no clue, ask @SolomonZelman

OpenStudy (jordanloveangel):

@SolomonZelman can u help me

OpenStudy (jordanloveangel):

@chucky00

OpenStudy (jordanloveangel):

@DullJackel09

OpenStudy (solomonzelman):

Lets say you have 2 points, (a,b) and (c,d) . The slope between them will be: \(\LARGE\color{black}{ slope=\frac{(b-d)}{(a-c)} }\) For example, if you have (-3,4) and (-2,-5) Then the slope will be \(\LARGE\color{black}{ slope=\frac{4-(-5)}{-3-(-2)} =\frac{4+5}{-3+2} =\frac{9}{-1}=-9 }\)

OpenStudy (solomonzelman):

(The y-values over the x-values ) So how would you do your problem, when you have (2,5) and (0,-4) .

OpenStudy (jordanloveangel):

so i then answer is -9

OpenStudy (jordanloveangel):

but i need a fracetion for this problem @SolomonZelman